OpenStudy (anonymous):

Alan bought a fencing foil for $56. When he went to the cash register, the total was $60.20. What was the sales tax rate?

OpenStudy (anonymous):

60.20 - 56 = 4.20 $4.20 tax was charged on $56.00. To find the decimal, divide 4.20/56 = 0.075 Now, convert the decimal to a percent. 0.075*100 = 7.5%

OpenStudy (anonymous):

Thank you!